Nancy drives in traffic for 2 and 1/2 hours. After traffic dies down she travels another 4 and 1/2 hrs. at 42 mph. If the entire trip was 371 miles, how fast was she driving in traffic?

Answer by stanbon(75887)

Nancy drives in traffic for 2 and 1/2 hours. After traffic dies down she travels another 4 and 1/2 hrs. at 42 mph. If the entire trip was 371 miles, how fast was she driving in traffic?
-------
1st leg DATA:
time = 5/2 hrs ; rate = x mph ; distance = (5/2)x miles
-------
2nd leg DATA:
time = 9/2 hrs ; rate = 42 mph ; distance = (9/2)*42 = 189 miles
-------
Equation:
distance + distance = 371
(5/2)x + 189 = 371
-----
(5/2)x = 182
----
x = (2/5)182
---
x = 72.8 mph(rate of the 1st leg)
=========================================
Cheers,
Stan H.